Determine the zeros of the function y = k/2(sin(x)); 0 ≤ x ≤ π


1
Expert's answer
2022-02-02T08:46:29-0500

Solution:

"y = \dfrac k2(\sin x); 0 ≤ x ≤ π"

Put y=0

"\Rightarrow \dfrac k2(\sin x)=0
\\\Rightarrow \sin x=0
\\\Rightarrow x=0,\,x=\pi"